# Prediction of flow and elastic stresses in a viscoelastic turbulent channel flow using convolutional neural networks

### v1

@@ 3 Results
3.1 V-predictions
@@
A qualitative prediction field for V-predictions is shown in figure 3 (corresponding to an instant in the test dataset where the input wall-shear rate is higher than the mean wall-shear rate). We observe the predicted velocity fields to be visually well correlated with the DNS fields at different target wall-normal locations. The linear correlation coefficient between the predicted and true streamwise-velocity fluctuation field exceeded 99% for predictions at , and gradually declining but remaining above 80% at . The RMS quantities of the streamwise velocity-fluctuation fields at are predicted with less than 3%, 6% and 15% error, respectively. With an increasing separation distance (wall-normal distance between the wall fields and the target velocity-fluctuation fields), the fields are less-correlated and thereby the performance of the network also decreases. Because of this, the RMS-normalized mean-absolute errors in the predicted streamwise-velocity fluctuations are 0.14, 0.29 and 0.47 at , respectively (see also figure 4).
@@
The MAE in the wall-normal and spanwise fluctuation fields remained below in the different target wall-normal locations considered in the study. However, the values in the wall-normal and spanwise velocity fluctuations are at least twice as large as those obtained in the RMS prediction of the streamwise component at the respective wall-normal locations. This is due to the influence of the polymers, which reduce turbulence by opposing the downwash and upwash flows generated by near-wall vortices (Dubief et al., 2004, 2005). Due to the absence of polymeric-stress information in the inputs to the network model for V-predictions, an accurate representation of the turbulence statistics in the spanwise and wall-normal fluctuation components becomes challenging.
@@
It should be emphasized that the network model is explicitly optimized for predicting instantaneous fields rather than reproducing the turbulence statistics. This emphasis is rooted in the motivation for non-intrusive sensing in an experimental setting, aimed at understanding the near-wall dynamics of viscoelastic turbulent channel flow. In addition, optimizing network models to accurately replicate turbulence statistics obtained from DNS could lead the model to learn the mean-flow behavior with a lower . This may also entail a compromise, as predictions during hibernating intervals could potentially become less accurate.
@@
When assessing the accuracy of the instantaneous predictions based on mean-absolute errors, as illustrated in figure 4, it becomes apparent that the MAE (in each test sample) varies with wall-shear rate for different target wall-normal locations. Specifically, in instances of low-wall-shear-rate, the absolute errors are notably lower, and increase with wall-shear rate. This is due to the fact that low-drag events typically exhibit reduced fluctuation intensity and it increases with wall-shear rate, leading to increased concentration of energy in small-scale features. Consequently, the network encounters relative difficulty in accurately predicting these small-scale features, resulting in higher prediction errors at large wall-shear-rate inputs. It is worth noting that the variation of MAE (in each test sample) with wall-shear rate stems from the selection of the loss function utilized in the network. Nevertheless, the obtained network model exhibits superior predictive performance in capturing velocity-fluctuation fields during low-wall-shear rate events. This observation underscores the potential utility of such models in obtaining sufficiently accurate velocity fluctuations in an experimental setting, more particularly for studying hibernation events in detail.
3.2 E-predictions, V-E-predictions
@@
A sample predicted field (corresponding to the same wall inputs as in figure 3) for E-predictions and V-E-predictions is shown in figure 5. Overall, the large-scale features in the polymer-stress quantities of interest are visually well-corresponding. For E-predictions, where polymeric stresses are predicted from DNS velocity fields at the same location there is no separation distance between the input and target fields and the linear correlation coefficient between the predicted and DNS polymer-shear stress, as well as with the trace of the polymer stress remained more than 90% for the different target wall-normal positions. Moreover, remained below 15% for E-predictions of and at different target .
@@
Note that in V-E-predictions the polymeric stresses are predicted directly from the wall inputs, without having access to the true velocity fields at the wall-normal location where those stresses are predicted. Instead, predicted auxiliary velocity fields at that location (together with wall inputs) are used to predict the polymeric stress fields. The obtained errors of around indicate that a small error in predicting velocity-fluctuation fields significantly impacts the errors in predicting the polymeric-stress fields, indicating that the auxiliary velocity fluctuation fields in V-E-predictions lack certain information that is connected to the polymeric activity in the small wavelengths. Nevertheless, the large-scale structures in the predicted polymeric-stress fields for V-E-predictions exhibit a qualitative agreement with the reference, as observed in figure 5.
@@
Examining the accuracy of instantaneous predictions based on mean-absolute errors, as depicted in figure 6 for E-predictions and V-E-predictions, reveals a similar trend in MAE (in each test sample) with respect to wall-shear rate as observed in V-predictions. Overall, the absolute errors increase with wall-shear rate. Further, the magnitude of such absolute errors in the field is nearly doubled for V-E-predictions (figure 6, ) compared to E-predictions (figure 6, ). Moreover, the MAE in predicting polymer-stress quantities of interest remains relatively constant across various target wall-normal positions with respect to the corresponding RMS quantities for both E-predictions and V-E-predictions.
The RMS of the trace of the polymer stress decreases as we move away from the wall for the considered locations and consequently the absolute errors in predicting decrease as the wall-normal distance increases. One plausible explanation for this phenomenon lies in the increased presence of energetic large-scale structures in the diagonal components of polymer stress tensor with respect to the wall-normal distance (see figure 8). Further, the trace of the polymer stress exhibits an increasing anti-correlation with the streamwise velocity fluctuation as the wall-normal position increases as observed from figure 7, which is related to the polymers extracting turbulent kinetic energy (Dubief et al., 2005).
3.3 Power-spectral density
@@
The distribution of energy in the predicted and DNS data across different scales are compared through the spectral analysis as illustrated in figure 8. The results show that the neural-network models successfully reproduce the energy content in the streamwise velocity component (denoted by ) at different wavelengths. However, for the wall-normal velocity fluctuations () and spanwise velocity fluctuations (), the network models exhibit limitations in reconstructing energy at the smallest wavelengths and specifically such errors in the smallest scales increase with increasing target wall-normal position.
@@
The power-spectral density obtained for the polymeric-shear stress (denoted by ) and the trace of polymer stress () are depicted in figure 8 for different wall-normal positions. We observe that the energetic structures correspond to wavelengths that are almost one order of magnitude smaller than those observed in the velocity fluctuations (refer to figure 8). This reveals that the polymer activity is predominantly concentrated in small-scale structures compared to the flow scales. Consequently, this suggests that the employed neural network model needs to reconstruct fine-scale polymer stress fields from coarse energy-containing features in the velocity fluctuations.
@@
For the E-predictions shown in figure 8, we observe the ability of the model to reconstruct the features containing energy at different wavelengths more accurately with minimal errors observed in the smallest scales. However in the case of V-E-predictions, where the wall inputs to the network feature large-scale energy-containing features (not shown here), the performance of the network is reduced in reconstructing the energy distribution of features at smaller scales, and rather the model tends to predict the large-scale features in the polymer-stress fields.

### v2

@@ 3 Results
The predictions of the trained network models are compared with the data obtained from DNS. The performance is assessed from a qualitative point of view and subsequently from a quantitative aspect, based on predictions of instantaneous fields, turbulent statistics and the two-dimensional power spectral density. Further, the importance of small-scale velocity fluctuations to successfully retrieve the small-scale polymer stress fluctuations is highlighted.
3.1 Prediction of velocity-fluctuations
@@
The predicted velocity-fluctuations at different target wall-normal positions using streamwise and spanwise wall-shear rate and wall pressure are qualitatively inspected. A sample prediction of the instantaneous velocity-fluctuation fields for the case of V-predictions is shown in figure 3 (corresponding to an instant in the test dataset where the input wall-shear rate is higher than the mean wall-shear rate). We note that the predicted velocity fields are visually well correlated with the DNS fields at different target wall-normal locations. The linear correlation coefficient between the predicted and true streamwise-velocity fluctuation fields exceeds 99% for predictions at , and gradually declines but remains above 80% at . The RMS quantities of the streamwise velocity-fluctuation fields at are predicted with less than 3%, 6% and 15% error, respectively. With an increasing separation distance (wall-normal distance between the wall fields and the target velocity-fluctuation fields), the fields are less-correlated and thereby the performance of the network also decreases. Because of this, the RMS-normalized mean-absolute errors in the predicted streamwise-velocity fluctuations are 0.14, 0.29 and 0.47 at , respectively (see also figure 4). The performance metrics for different network models are summarized in Appendix A.
@@
The MAE in the wall-normal and spanwise fluctuation fields remained below in the different target wall-normal locations considered in the study (refer figure 4 for normalized quantities). However, the values in the wall-normal and spanwise-velocity fluctuations are at least twice as large as those obtained in the RMS prediction of the streamwise component at the respective wall-normal locations. This is due to the influence of the polymers, which reduce turbulence by opposing the downwash and upwash flows generated by near-wall vortices (Dubief et al., 2004, 2005). Due to the absence of polymeric-stress information in the inputs to the network model for V-predictions, an accurate representation of the turbulence statistics in the spanwise and wall-normal fluctuation components becomes challenging.
@@
It should be emphasized that the network model is explicitly optimized for predicting instantaneous fields rather than reproducing the turbulence statistics. This emphasis is rooted in the motivation for non-intrusive sensing in an experimental setting, aimed at understanding the near-wall dynamics of viscoelastic turbulent channel flow. In addition, optimizing network models to accurately replicate turbulence statistics obtained from DNS could lead the model to learn the mean-flow behaviour with a lower . This may also entail a compromise, as predictions during hibernating intervals could potentially become less accurate.
@@
When assessing the accuracy of the instantaneous predictions based on mean-absolute errors, as illustrated in figure 4, it becomes apparent that the MAE (in each test sample) varies with wall-shear rate for different target wall-normal locations. Specifically, in instances of low-wall-shear-rate, the absolute errors are notably lower and increase with wall-shear rate. This is due to the fact that low-drag events typically exhibit reduced fluctuation intensity, which increases with wall-shear rate, leading to increased concentration of energy in small-scale features. Consequently, the network encounters relative difficulty in accurately predicting these small-scale features, resulting in higher prediction errors at large wall-shear-rate inputs. It is worth noting that the variation of MAE (in each test sample) with wall-shear rate stems from the selection of the loss function utilized in the network. Nevertheless, the obtained network model exhibits superior predictive performance in capturing velocity-fluctuation fields during low-wall-shear rate events. This observation underscores the potential utility of such models in obtaining sufficiently accurate velocity-fluctuations in an experimental setting, more particularly for studying hibernation events in detail.
3.2 Prediction of polymer stress
@@
The predicted velocity fluctuations at various target wall-normal positions show a good agreement with the reference DNS quantities. Following this, the focus is to estimate the polymer stress quantities of interest based on velocity fluctuations. Additionally, an effort is made to retrieve polymer stress using only wall information, which is advantageous in experimental settings. This section begins with a qualitative analysis of the predicted polymer stress fields. A sample predicted field of polymer stress components (corresponding to the same time instant as the wall inputs shown in figure 3) for E-predictions and V-E-predictions is shown in figure 5. Overall, the large-scale features in the polymer-stress quantities of interest are visually in good agreement with DNS. For E-predictions, where polymeric-stresses are predicted from DNS velocity fields at the same location there is no separation distance between the input and target fields and the linear correlation coefficient between the predicted and DNS polymer-shear stress, as well as with the trace of the polymer stress remained more than 90% for the different target wall-normal positions. Moreover, remained below 15% for E-predictions of and at different target wall-normal distances.
@@
Note that in V-E-predictions the polymeric-stresses are predicted directly from the wall inputs, without having access to the true velocity fields at the wall-normal location where those stresses are predicted. Instead, predicted auxiliary velocity fields at that location (together with wall inputs) are used to predict the polymeric-stress fields. The obtained errors of around indicate that a small error in predicting velocity-fluctuation fields significantly impacts the errors in predicting the polymeric-stress fields (see also §3.4), suggesting that the auxiliary velocity-fluctuation fields in V-E-predictions lack certain information that is connected to the polymeric activity in the small wavelengths. Nevertheless, the large-scale structures in the predicted polymeric-stress fields for V-E-predictions exhibit a qualitative agreement with the reference, as observed in figure 5.
@@
Examining the accuracy of the instantaneous predictions based on mean-absolute errors, as depicted in figure 6 for E-predictions and V-E-predictions, reveals a similar trend in MAE (in each test sample) with respect to wall-shear rate as observed in V-predictions. Overall, the absolute errors increase with wall-shear rate. Further, the magnitude of such absolute errors in the field is nearly doubled for V-E-predictions (figure 6, ) compared to E-predictions (figure 6, ). Moreover, the MAE in predicting polymer-stress quantities of interest remains relatively constant across various target wall-normal positions with respect to the corresponding RMS quantities for E-predictions and with marginal distinctions in V-E-predictions.
3.3 Energy distribution across different length scales
The statistical analysis of the predicted velocity fluctuations using V-predictions and the predicted polymer stresses using E-predictions closely matches the results obtained from DNS. However, there is a significant deviation in the second-order statistics of polymer stress quantities predicted using only the wall-shear rate and wall pressure (V-E predictions). To understand the reasons behind this, the distribution of energy across different length scales in the FCN predictions is examined and compared with the reference DNS.
@@ 3.3 Energy distribution across different length scales
The distribution of energy in the predicted and DNS data across different scales are compared through spectral analysis as illustrated in figure 7. The results show that the neural-network models successfully reproduce the energy content in the streamwise velocity component (denoted by ) at different wavelengths. However, for the wall-normal velocity fluctuations () and spanwise velocity fluctuations (), the network models exhibit limitations in reconstructing energy at the smallest wavelengths and specifically such errors in the smallest scales increase with increasing target wall-normal position.
@@
The power-spectral density obtained for the polymeric-shear stress (denoted by ) and the trace of polymer stress () are depicted in figure 7 for different wall-normal positions. We observe that the energetic structures correspond to wavelengths that are almost one order of magnitude smaller than those observed in the velocity fluctuations (refer to figure 7). This reveals that the polymer activity is predominantly concentrated in small-scale structures compared to the flow scales. Consequently, this suggests that the employed neural-network model needs to reconstruct fine-scale polymer stress fields from coarse energy-containing features in the velocity-fluctuations.
@@
For the E-predictions (figure 7), where exact velocity fields from DNS at the same wall-normal position are used as inputs, we observe the ability of the model to reconstruct the features containing energy at different wavelengths more accurately with minimal errors observed in the smallest scales. However in the case of V-E-predictions, where the wall inputs to the network feature large-scale energy-containing features (not shown here), the performance of the network is reduced in reconstructing the energy distribution of features at smaller scales, and rather the model tends to predict the large-scale features in the polymer-stress fields. These observations underscore the importance of providing accurate velocity-fluctuation fields as inputs to the FCN. Although the auxiliary velocity fields have a very small error compared to DNS velocity fields (give a number), they lack certain information when reconstructed by FCN using wall inputs in V-E-predictions, the network is unable to capture the polymer-stress features at small and medium scales.
3.4 Effects of small-scale velocity fluctuations on the polymer stress predictions
Spectral analysis reveals that the energetic scales in velocity fluctuations are approximately one order of magnitude larger than those in the polymer stress quantities of interest. Consequently, it becomes challenging for the FCN to reconstruct polymer stress information at finer scales using the large energy-containing structures in velocity fluctuations. Furthermore, for E-predictions, DNS velocity fluctuation fields are used as inputs whereas for V-E-predictions, the (auxiliary) inputs are the predicted velocity fluctuations at a target wall-normal position that exhibits a good agreement with the DNS velocity fluctuation fields. It should be highlighted that the (auxiliary) velocity fluctuation fields employed in V-E-predictions quite accurately comprise the large-scale features as observed with DNS fields, although they exhibit certain differences in the small-energy containing features (see also figure 7). Therefore, if the large energy-containing structures in velocity fluctuations are responsible for the accurate prediction of polymer stresses, there should be minimal differences between E and V-E predictions of polymer stress components. However, we observe that small-scale polymer stress fluctuations in V-E predictions are under-predicted compared to E-predictions (see figure 5). This indicates that small-scale velocity fluctuations in DNS fields are crucial for the accurate prediction of polymer stresses. Thereby, the observations also signify that polymer stresses and velocity fluctuations are strongly coupled at fine, low-energetic scales.
We investigate the relationship between polymer stress and velocity fluctuations at smaller scales by systematically removing the turbulent kinetic energy at these finer scales and observing the resulting polymer stress predictions by FCN. The low-pass filtering of velocity fluctuations (as outlined in §2.3) is performed to evaluate the accuracy of prediction of polymer stresses by utilizing only the large energy-containing features in velocity fluctuations. This approach not only helps in identifying the effects of small-scale features in the input velocity-fluctuation fields on the prediction of polymer stress but also enables a scale requirement recommendation for velocity fluctuations from potential experimental investigations to estimate the polymer stress information. The effect of filtering velocity-fluctuations using the pre-multiplied two-dimensional (2D) power-spectral density (PSD) for the velocity-fluctuations () at as shown in figure 8. The figure illustrates the effect of filtering the small-scale velocity-fluctuations with a wavelength threshold of .
A sample instantaneous field of the trace of polymer stress in the test dataset is shown in figure 9. The reference DNS field at different wall-normal locations is shown alongside the corresponding predictions obtained from FCN, with respective inputs at different cut-off wavelengths. It is observed that the FCN predictions successfully capture all the different scales present in the reference DNS. As the cut-off wavelength in the input velocity-fluctuations to FCN increases, the predicted outputs begin to lack certain small-scale features. At , where the absence of small scales contributes to a relative loss () of less than of turbulent kinetic energy in the inputs at different considered here, the errors in the prediction of small scale become significant, as illustrated in figure 9. Further for , constituting a relative loss of less than of TKE of inputs, we observe that the small-scale features are increasingly depleted in the predictions. Nevertheless, the predictions still manage to capture the regions with large polymer extension, where is higher. A similar observation for the polymer shear-stress is provided in Appendix B.
Overall, the variation of errors in predicting the corresponding RMS quantities of the polymer stress quantities of interest is shown in figure 10 for different values of cut-off wavelengths of inputs. The corresponding normalized mean-absolute errors are also indicated. The shaded region in the plot corresponds to the input velocity-fluctuations provided to the FCN, corresponding to a relative loss () of less than of turbulent kinetic energy due to filtering. Overall, the mean errors observed at different target wall-normal locations exhibit a logarithmic growth with respect to the cut-off wavelength of the input velocity-fluctuations. This also explains the observation of higher errors in V-E predictions when using wall inputs to predict polymer stress quantities of interest at different compared to E-predictions. Although the predicted velocity-fluctuations from V-predictions resemble close to that of DNS reference, there is a logarithmic increase in prediction errors of polymer stresses when utilizing the auxiliary predictions of velocity-fluctuations which lacks energy content at certain scales (see figure 7).
In addition, for retrieval of polymer-stress information from possible near-wall experimental velocity fields using FCN, it becomes necessary to resolve finer scales, typically lower than 10 viscous lengths, to obtain more reliable polymer stress behaviour at the smallest scales. This hypothesis will be confirmed more conclusively in the next section where we further examine the effects of artificially excluding small scales from the DNS velocity fields.
3.5 Spectral analysis on the effects of small-scale velocity fluctuations
Identifying the influence of small-scale features in the velocity-fluctuation fields on the polymer stress, we probe the distribution of energy in the predicted polymer stresses due to the absence of small, low-energetic scales in the velocity fluctuations. The pre-multiplied two-dimensional power-spectral density of the trace of polymer stress ( at is depicted in figure 11. Figure 11 shows the 10, 50 and 80 contours of reference (DNS) polymer stress by shaded contours, against the FCN predictions of trace of polymer stress using DNS velocity-fluctuation fields as input to FCN () (black contour lines) and filtered velocity fluctuations as input to FCN () (purple contour lines).
The predictions using exact velocity fields closely capture the energy content across different scales compared to the reference (DNS) polymer stress (black contours are almost identical to shaded contours in figure 11). However, it should be noted that there are always some small errors in predicting the trace of polymer stress fields using DNS velocity-fluctuations as input. These errors stem from the absence of certain correlations in the inputs, such as the wall-normal gradients of velocity-fluctuations and the time history of velocity-fluctuations.
On the other hand, filtering the input velocity-fluctuations with so that they lack the very finest scales, results in a very significant deviation in the power density spectra of the predicted polymer stress (purple contours are very different from shaded contours). Most notably, the energy-containing structures in the predictions are at larger spatial scales than in the DNS. Additionally, the predictions now lack certain small-scale energy components. This shows that polymers interact with very small scales in the velocity fields in viscoelastic turbulence. The similarity between purple contours (predictions with filtered velocity fields) and green contours in figure 7 (V-E predictions) confirms that errors in V-E predictions stem from the absence of the smallest scales in the auxiliary velocity fields.
The difference between the FCN predictions and the DNS reference polymer stress field represents the error in the predictions. To analyse this error quantitatively, we show the error spectra in figure 11. The error spectra illustrate the difference between the reference and predicted polymer-stress-fluctuations using DNS velocity-fluctuations as input (-, black dashed contour), and the difference between reference (DNS) and predicted polymer-stress-fluctuations using filtered velocity-fluctuations with a cut-off wavelength of , as input (-)) is depicted by a purple dashed contour). From figure 11, we observe that the errors are highest at the small yet energetic scales of the trace of polymer stress, while large scales are less affected by errors. This confirms the visual finding from figure 5, that the largest scales are relatively well captured by V-E predictions.
Finally from figure 11, we examine the effect of filtering small scales in the velocity-fluctuations on the differences introduced in the predictions by the FCN. Here, the spectra of the difference between the predicted polymer-stress-fluctuations with DNS inputs (refer to figure 11, black contour lines) and the predicted polymer-stress-fluctuations (refer to figure 11, purple contour lines) with filtered inputs (-) is plotted with brown contour lines. In addition, we include the spectra of the difference between reference DNS velocity-fluctuations and filtered velocity-fluctuations with for components (respectively indicated with blue, orange and green contour lines). The latter contours are very narrow, which shows that changes in velocity fields due to filtering mostly occur at a narrow range of scales. However, predictions of polymer stresses are affected at a strikingly large range of scales (brown contours). Figure 11 hence shows that the loss of turbulent kinetic energy in the small-scales in the velocity-fluctuations does not necessarily correspond to the loss of energy at similar scales in the polymer stress. This observation is aligned with the studies of Nguyen et al. (2016), where they showed from scale-to-scale analysis (Casciola & de Angelis, 2007; Xi et al., 2013; Valente et al., 2014) of isotropic viscoelastic turbulence, that the loss of turbulent kinetic energy at a given small scale does not necessarily correspond to a gain of polymer energy at the same scale, and vice versa.
3.6 Interpretation of predictions with filtered velocity fluctuations
Overall, we observe that the small-scale features in the velocity-fluctuation fields are crucial for the estimation of polymer stress. In this section, the reason for such an observation is hypothesized based on polymer-flow interaction and obtained predictions from FCN. We start with the observation of the probability distribution function of polymer stress. The probability distribution function for the trace of polymer stress is shown in figure 12. The predictions from FCN using DNS input velocity-fluctuations closely capture the resulting probability distribution function observed with the DNS stress fields in the test dataset. However, utilizing the filtered velocity-fluctuation fields with results in the distribution of the data to be closer to the mean value. The negative and positive fluctuations of the trace of the polymer stress are under-predicted indicating that the relaxation behaviour of the polymer is over-predicted and the maximum stretch of polymers is under-predicted by FCN. The over-prediction and under-prediction of polymer stress may stem from the fact that the FCN is optimized to capture the mean behaviour of the polymer-stress-fluctuations for the given velocity-fluctuation input. Because of the alteration of the input velocity-fluctuation field via filtering of small scales, the FCN is unable to predict well the extreme behaviour of polymer stress and thereby converges close to the mean behaviour and hence the resulting prediction of the standard deviation of the polymer stress quantities of interest are lower in comparison to that observed with the DNS fields. Consequently, this contributes to growing errors with respect to the cut-off wavelength of input velocity-fluctuations. However, there may be physical reasons attributed to the difficulty in predicting polymer stress without small-scale features in velocity fluctuations that are independent of the chosen technique to model the relationship between polymer stress and velocity fluctuations. Future work will involve performing direct numerical simulations with filtered velocity fluctuations to identify the effects of small-scale velocity fluctuations on the polymer stress fluctuation fields. Additional tasks also include transfer learning based approaches to evaluate the performance of the network models at different Weissenberg numbers.
Filtering small scales in the velocity-fluctuation fields results in an alteration of the turbulent kinetic energy of the flow. In the buffer region that is considered in the study, there exists an anti-correlation between streamwise velocity-fluctuation and the fluctuations of the trace of polymer stress (as shown in figure 13), thereby an increase in polymer stresses (and consequently polymer energy) leads to decrease in the turbulent kinetic energy of the fluid (see also figure 14).
The RMS-normalized joint probability distribution of the prediction errors in the trace of polymer stress with respect to turbulent kinetic energy is probed in figure 14. From the figure, the relationship between the trace of polymer stress (indicative of polymer stretch and elastic energy stored by polymers) and the turbulent kinetic energy of the flow from DNS samples in the test dataset is observed. It is evident that the polymers are highly stretched in regions with low turbulent kinetic energy and vice versa. This suggests that polymers extract turbulent kinetic energy from the fluid, particularly in the considered wall-normal position in the buffer region. Moreover, the relaxation of the polymer stress, which can occur in the core of the streamwise vortices, releases energy into the fluid flow (Xi, 2019).
From figure 14, it is observed that the under-prediction of the trace of polymer stress fluctuations is significant in the regions of lower turbulent kinetic energy. This indicates that the effect of filtering small scales results in over-prediction of relaxation behaviour of polymer stress and under-prediction of maximum polymer elongation and thereby maximum trace of polymer stress, particularly in the regions with lower turbulent kinetic energy. Thus, due to the alteration of the distribution of small-scale input velocity-fluctuations, the FCN model is unable to reproduce the features of energy extraction by polymers from the turbulent flow. Hence, it is pertinent to capture accurately the turbulent kinetic energy in the flow and that small-scale velocity-fluctuations are required to accurately capture the extraction of turbulent kinetic energy from flow by polymers.
The results from the present study suggest that polymer stress can be estimated by directly providing potential experimental measurements of wall quantities (V-E predictions) or near-wall velocity fluctuations (E predictions) as inputs to the network model. It is important to note that experimental fields may contain noise, necessitating retraining of the network model using transfer learning methodologies to optimize the model weights with the acquired experimental dataset. However, in the current supervised learning framework, such approaches require the presence of reference experimental stress fields for retraining the network models, which is not applicable in this case. It was recently shown that Lagrangian stretching fields extracted from particle image velocimetry (PIV) serve as indicators of polymer elongation fields under certain conditions (Kumar et al., 2023), but their applicability to stress field substitutes in channel flow turbulence remains to be investigated. Therefore, we propose that as an initial step to gain insight into elastic stresses from experimental measurements, interpolation (for matching resolution) and de-noising techniques (Raiola et al., 2015; Nekkanti & Schmidt, 2021; Yousif et al., 2023c) can be employed to retrieve accurate wall fields or velocity fluctuations from experimental data. These refined velocity inputs can then be coupled with the proposed DNS-trained network models to obtain predictions of instantaneous elastic stress fields within the probed domain.
The present work can also be extended towards identifying coherent structures in the flow and thereby employing SHapley Additive exPlanations (SHAP) algorithm to explain the importance of the coherent structure (Cremades et al., 2024) which can improve the understanding of the dynamical role of coherent structures in viscoelastic turbulent flow. Hence the present work serves as a starting point, holding potential for many future works to understand better the polymer physics in drag reduced flows.